OpenStudy (anonymous):

Find the LCM. t^3+8t@+16t, t^2-8t

jhonyy9 (jhonyy9):

what is there 8t@ ?

OpenStudy (turingtest):

t^3+8t^2+16t,t^2-8t t^3+8t^2+16t=t(t+4)^2 t^2-8t=t(t-8) so the factors of the LCM are t(t-8)(t+4)^2 BTW I already solved this for you

OpenStudy (anonymous):

Sorry its 8t^2

OpenStudy (turingtest):

that is what I solved above

OpenStudy (anonymous):

yeah thx it was sapoto look like t(t+4)^2(t-8)